OpenStudy (anonymous):

If tan A - tan B = x and cot B - cot A =Y , pt COT (a-b) = 1/X + 1/Y

OpenStudy (anonymous):

@akash123

OpenStudy (anonymous):

sin(A-B)/ cosA*cosB =1/x from Tan A - tan B=x

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

and sin(A-B)/ sinA* SinB=y from cotB- cotA =y sin(A-B)/ cosA*cosB =x from Tan A - tan B=x tanA*tanB= x/y

OpenStudy (anonymous):

@akash123 sin(A-B)/ cosA*cosB =x from Tan A - tan B=x

OpenStudy (anonymous):

cot(A-B) = [ cot A* cot B+1] / ( cot B - cot A)

OpenStudy (anonymous):

yes...thanks...I have corrected it later

OpenStudy (anonymous):

sin(A-B)/ sinA* SinB=y from cotB- cotA =y sin(A-B)/ cosA*cosB =x from Tan A - tan B=x divide 2nd by 1st then u'll get tanA*tanB= x/y

OpenStudy (anonymous):

yes i understood that !

OpenStudy (anonymous):

now use the identity cot(A-B) = [ cot A* cot B+1] / ( cot B - cot A) and substitute the value of cot A* cot B and (cot B- cot A) then u'll get the answer

OpenStudy (anonymous):

another way @sauravshakya write the reciprocal f 1st and 2nd equation and add both then simplify it.

OpenStudy (anonymous):

Yep... I would go with that....

OpenStudy (anonymous):

okay got it :) thank you !

OpenStudy (anonymous):

:)